I can't show what the power is so i'll just say to the ___ power.
3 to the x+2 power = 5 to the x-1 power. I'm not sure how to do it when there are x's on both sides and they normally cancel out in previous problems.

Answer by stanbon(75887)

3 to the x+2 power = 5 to the x-1 power.
---
3^(x+2) = 5^(x-1)
---
(x+2)log(3) = (x-1)log(5)
---
[log(3) - log(5)]x = -[2log(3) + log(5)]
Do the arithmetic and divide to get:
x = 7.452..
Cheers,
Stan H.